OpenStudy (anonymous):

6x - 8 + 2x - 5 = 7 - 2x x =

OpenStudy (imtant):

Bring all x to the left, 6x+2x+2x=8+5+7 So 10x=20 x=?

OpenStudy (anonymous):

x = 10

OpenStudy (imtant):

Yes,correct.

OpenStudy (anonymous):

thank you bby

OpenStudy (imtant):

Wc..